A portfolio has expected return of 13.2 percent and standard deviation of 18.9 percent. Assuming that the returns of the portfolio are normally distributed, what is the probability that, in any given year, the return of the portfolio will be greater than -5.7 percent.

A.

0.950

B.

0.840

C.

0.680

D.

0.975
Answer #1
as per 68% rule
range = mean +/- 1*std dev
=13.2 +/- 1*18.9
=13.2 - 1*18.9 to 13.2 + 1*18.9
=-5.7 to 32.1

As lower bound of 68% rule = -5.7 probability of having value above -5.7 = 0.68/2+0.5=0.84